# Context-Free Synthetic Data Mitigates Forgetting

### v1

@@ 3.1 Pretrained-only models
(A) We consider Olmo-1B Groeneveld et al. (2024). Olmo has publicly available pretraining data Soldaini et al. (2024), and is known to be not pretrained on math Groeneveld et al. (2024); hence it’s 0-shot GSM8K numbers are bad.
@@
(C) We train Olmo on MetaMath-QA Yu et al. (2023), and evaluate the downstream improvement by GSM8K. Fine-tuning on this dataset helps the model improve on this benchmark, but leads to forgetting of its pre-existing performance; this sets the stage for evaluating our (and other) methods.
Pre-existing Tasks
Commonsense MMLU BB Hard AGIEval Avg. GSM8K
Base 50.35 24.36 25.56 17.73 29.50 1.59
FT 40.89 23.06 11.49 18.77 23.55 29.49
CS 42.97 24.16 14.77 18.28 25.05 19.71
P 46.67 23.01 10.69 18.56 24.73 15.09
CFS 47.40 23.38 19.37 19.37 27.38 26.00
Table 2: This table studies the pretrained-only (i.e. Olmo-1B, MetaMathQA, GSM8K) setup described in Section 3.1. We see that standard finetuning FT is better than on the downsream task (GSM8K) but worse on the pre-existing tasks; this is forgetting. The remaining three rows show that our method CFS is much more effective at mitigating forgetting as compared to the other two data-augmentation-based methods P and CS.
Comparison against data-augmentation methods In Table 2, we see that standard finetuning of Olmo on MetaMathQA leads to big increase in it’s GSM8K performance, though it’s pretraining abilities go down - this is forgetting. Including contextually generated data into the mix – i.e. CS – hurts GSM8K evaluation. This is intuitive as augmenting with contextually generated data brings incorrect solutions to MetaMathQA questions into the mix. Context-free generation help the model to retain it’s pretraining abilities and learn the downstream task. Surprisingly context-free generations are better than including model’s pretraining data into the mix.
Comparing against weight-space methods In Table 3 we compare our method with relevant baselines. The main results are encapsulated in the table’s caption. Importantly we see that we are better than the relevant baselines. For $l_{2}$ regularization we take the the regularization penalty as $31e-$ 3. We do a through sweep over the hyper-parameters for baselines and report the relevant numbers here. Specifically we sweep over the ranks of the low-rank matrices in LoRA (Table 8), over regularization penalty for $l_{2}$ (Table 10) and over averaging rations $\alpha$ for Wise-FT (Table 11. CFS provides a better trade-off than these methods.
Pre-existing Tasks
Commonsense MMLU BB Hard AGIEval Avg. GSM8K
Base 50.35 24.36 25.56 17.73 29.50 1.59
FT 40.89 23.06 11.49 18.77 23.55 29.49
LoRA $(r=256)$ 48.41 23.71 21.87 18.72 28.18 20.17
$\ell_{2}$ regularization 47.43 23.03 18.92 19.76 27.28 28.43
Wise-FT$(\alpha=0.5)$ 45.53 23.10 19.64 18.85 26.78 14.56
CFS (50%) 46.96 23.29 23.19 19.52 28.24 29.34
Table 3: This table studies the pretrained-only (i.e. Olmo-1B, MetaMathQA, GSM8K) setup described in Section 3.1. Here we compare our data-augmentation based method CFS to the popular weight-space approaches LoRA, $\ell_{2}$ and Wise-FT. Each of these three baselines have hyper-parameters which we optimized over (see appendix); we report their best numbers here. We similarly ablate over the amount of context-free samples in our method (and find 50% – i.e. half as many synthetic as the number in the finetuning dataset to be optimal). While all these methods mitigate forgetting, our CFS seems the most effective.
Ablations We study our methods performance on ablating the sampling temperature and number of samples we generate for augmentations.
Model Avg. GSM8K
Base 29.50 1.59
FT 23.55 29.49
$T=0.6$ 27.38 25.47
$T=0.8$ 26.89 26.31
$T=1.0$ 27.40 26.00
$T=1.2$ 26.98 26.54
Model Avg. GSM8K
Base 29.50 1.59
FT 23.55 29.49
10% 27.59 28.43
50% 28.24 29.34
100% 27.38 26.00
200% 27.81 22.52
@@
Table 4: In these tables, we ablate over sampling temperature (Left) and number of generated synthetic samples (Right) for the pretrain-only setup of Section 3.1. Each table reports the model’s average pre-existing tasks performance as Avg. and it’s GSM8K accuracy. We see that our method is robust to choice of temperature (Left) used to sample, attenuating forgetting nonetheless. For the number of samples table (Right), different rows here correspond to different percentages (compared to downstream dataset size) of generated samples. For e.g., for 10% we generate a total of 10% of MetaMathQA dataset size = 40K samples for augmentation. FT can also be considered to be 0%. We see just 10% samples are enough for attenuating forgetting.
Main results are encapsulated in the caption of Table 4. We see that our method is robust to choice of temperature used to sample, attenuating forgetting nonetheless. We choose $T=1$ for rest of our experiments. The number of synthetic data samples used for augmentations display the expected trend with the pre-existing task average increasing and the GSM8K performance decreasing as we increase the augmentation data. We fix the number of samples to be 100%, i.e., equal to dataset size for our experiments.
3.2 Reasoning models
@@
(A) We consider R1-Distill-Llama-8B DeepSeek-AI (2025). R1-Distill-Llama-8B is made by finetuning Llama-3.1-8B-Instruct et al. (2024) on R1 reasoning traces and is a state-of-the-art 8B reasoning model.
@@
(C) We finetune R1-Distill-Llama-8B on the MedReason dataset Wu et al. (2025) following the setup of Wu et al. (2025). We evaluate the model’s medical abilities by MedQA Jin et al. (2021), MedBulletsOp4 (MBOp4), MedBulletsOp5 (MBOp5) Wu et al. (2025). MedReason dataset has 32K samples and is constructed by converting medical question-answer pairs into reasoning steps grounded in medical knowledge-graph. This is done by structured prompting of state-of-the-art LLMs [^3] MedReason huggingface link : https://huggingface.co/datasets/UCSC-VLAA/MedReason.
@@ Results
We also lack a comparison with $\ell_{2}$ regularization for this setup. This is because $\ell_{2}$ requires loading two models onto the GPU (the initial model and the fine-tuned model) while training. For making training these large models feasible we use DeepSpeed Zero3 parameter partition which we found is not amiable to accessing the target model parameters while training. We defer this and is on our future work.
@@ 4 Limitations
Appendix
@@ Appendix C Evaluation details
  • Commonsense reasoning datasets : Following Groeneveld et al. [2024] we average the performance on the following datasets for commonsense reasoning : 1. ARC-challenge, 2. ARC-easy, 3. Boolq, 4. Hellaswag, 5. Openbookqa, 6. Piqa, 7. Siqa, 9. Winogrande
  • We borrow the setup from et al. [2024] and report numbers considered in their "general language tasks" table for pretrained models. Since Olmo is not pretrained on math and code we don’t report those numbers and we found reading comprehension numbers for the models to be unreliable.
@@
  • We use the standard sampling params (from  DeepSeek-AI [2025]) for evaluating reasoning performance : temperature = 0.6, and probability threshold of 0.95. We generate for max length of 32768. We use vLLM for generating responses and inherit the reproducibility issues in it’s generation. We also append a <think> token to the query whenever we are generating a response.
@@ E.2 L2 Complete Table
See Table 10 for $\ell_{2}$ hyperparameter sweep for Olmo. $11e-$ 1 proves to be too strong of a regularization, we take $31e-$ 3 as having the best of both worlds.

### v2

@@ 3.1 Pretrained-only models
(A) We consider Olmo-1B (Groeneveld et al., 2024). Olmo has publicly available pretraining data (Soldaini et al., 2024), and is known to be not pretrained on math (Groeneveld et al., 2024); hence it’s 0-shot GSM8K numbers are bad.
@@
(C) We train Olmo on MetaMath-QA (Yu et al., 2023), and evaluate the downstream improvement by GSM8K. Fine-tuning on this dataset helps the model improve on this benchmark, but leads to forgetting of its pre-existing performance; this sets the stage for evaluating our (and other) methods.
Sampling Temp. Number of Generations
Model Avg. GSM Model Avg. GSM
Base 29.50 1.59 Base 29.50 1.59
FT 23.55 29.49 FT 23.55 29.49
$0.6$ 27.38 25.47 10% 27.59 28.43
$T=0.8$ 26.89 26.31 50% 28.24 29.34
$T=1.0$ 27.40 26.00 100% 27.38 26.00
$T=1.2$ 26.98 26.54 200% 27.81 22.52
@@
Table 3: In these tables, we ablate over sampling temperature (Left) and number of generated synthetic samples (Right) for the pretrain-only setup of Section 3.1. Each table reports the model’s average pre-existing tasks performance as Avg. and it’s GSM8K accuracy. We see that our method is robust to choice of temperature (Left) used to sample, attenuating forgetting nonetheless. For the number of samples table (Right), different rows here correspond to different percentages (compared to downstream dataset size) of generated samples. FT can also be considered to be 0%. We see just 10% samples are enough for attenuating forgetting.
@@
Comparison against data-augmentation methods In Table 2, we see that standard finetuning of Olmo on MetaMathQA leads to big increase in it’s GSM8K performance, though it’s pretraining abilities go down - this is forgetting. Including contextually generated data into the mix – i.e. CS – hurts GSM8K evaluation. This is intuitive as augmenting with contextually generated data brings incorrect solutions to MetaMathQA questions into the mix. Context-free generation help the model to retain it’s pretraining abilities and learn the downstream task. Surprisingly context-free generations are better than including model’s pretraining data into the mix. Note that Olmo-1B is a pretrained-only model. Hence, mixing with pretraining data is equivalent to data-replay methods in continual learning. Our results suggest that in context of LLMs, context-free generations can outperform even data-replay. We hypothesize that the high variance and the noisy inherent in the pretraining data might be be the reason for the sub-optimal performance.
Comparing against weight-space methods In Table 2 we compare our method with relevant baselines. The main results are encapsulated in the table’s caption. Importantly we see that we are better than the relevant baselines. For $l_{2}$ regularization we take the the regularization penalty as 1e31e-3. We do a through sweep over the hyper-parameters for baselines and report the relevant numbers here. Specifically we sweep over the ranks of the low-rank matrices in LoRA (Table 7), over regularization penalty for $l_{2}$ (Table 9) and over averaging rations $\alpha$ for Wise-FT (Table 10. CFS provides a better trade-off than these methods.
Ablating on number of generated samples We study our methods performance on ablating the number of samples we generate for augmentations. Main results are encapsulated in the caption of Table 3. The number of synthetic data samples used for augmentations display the expected trend with the pre-existing task average increasing and the GSM8K performance decreasing as we increase the augmentation data. We would like to highlight that in the regime when the generation budget is limited (i.e., just 10% of the size of the finetuning data), CFS is still able to effectively mitigate forgetting.
Ablating on generation parameters We study our methods performance on ablating the sampling temperature in Table 3. We see that our method is robust to choice of temperature used to sample, attenuating forgetting nonetheless. We choose $T=1$ for rest of our experiments. We skip the ablations on other generation parameters like nucleus sampling parameter and the CFS generation prompt.
3.2 Reasoning models
@@
(A) We consider R1-Distill-Llama-8B (DeepSeek-AI, 2025). R1-Distill-Llama-8B is made by finetuning Llama-3.1-8B-Instruct (et al., 2024) on R1 reasoning traces and is a state-of-the-art 8B reasoning model.
@@
(C) We finetune R1-Distill-Llama-8B on the MedReason dataset (Wu et al., 2025) following the setup of (Wu et al., 2025). We evaluate the model’s medical abilities by MedQA (Jin et al., 2021), MedBulletsOp4 (MBOp4), MedBulletsOp5 (MBOp5) (Wu et al., 2025). MedReason dataset has 32K samples and is constructed by converting medical question-answer pairs into reasoning steps grounded in medical knowledge-graph. This is done by structured prompting of state-of-the-art LLMs [^3] MedReason huggingface link : https://huggingface.co/datasets/UCSC-VLAA/MedReason.
@@ Results
We lack a comparison with $\ell_{2}$ regularization for this setup. This is because $\ell_{2}$ requires loading two models onto the GPU (the initial model and the fine-tuned model) while training. For making training these large models feasible we use DeepSpeed Zero3 parameter partition which we found is not amiable to accessing the target model parameters while training. We defer this and is on our future work.
@@ 4 Limitations
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix C Evaluation details
  • Commonsense reasoning datasets : Following (Groeneveld et al., 2024) we average the performance on the following datasets for commonsense reasoning : 1. ARC-challenge, 2. ARC-easy, 3. Boolq, 4. Hellaswag, 5. Openbookqa, 6. Piqa, 7. Siqa, 9. Winogrande
  • We borrow the setup from (et al., 2024) and report numbers considered in their "general language tasks" table for pretrained models. Since Olmo is not pretrained on math and code we don’t report those numbers and we found reading comprehension numbers for the models to be unreliable.
@@
  • We use the standard sampling params (from  (DeepSeek-AI, 2025)) for evaluating reasoning performance : temperature = 0.6, and probability threshold of 0.95. We generate for max length of 32768. We use vLLM for generating responses and inherit the reproducibility issues in it’s generation. We also append a <think> token to the query whenever we are generating a response.
@@ E.2 L2 Complete Table
See Table 9 for $\ell_{2}$ hyperparameter sweep for Olmo. 1e11e-1 proves to be too strong of a regularization, we take 1e31e-3 as having the best of both worlds.